OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7.

OpenStudy (anonymous):

first sub that point into f(x) = mx + b along with the same slope of that second equation and solve for b an equation that is parallel has the same slope but a different y -int

OpenStudy (anonymous):

Okay well ive done some research online and ive reviewed several pages on how to solve this, I came up with y = 4x -2 but thats not one of my answers :(

OpenStudy (anonymous):

what are your answers ?

OpenStudy (anonymous):

y = 4x + 10 y = 1/4x + 5/2 y = -4x - 6 y = -1/4x + 3/2

OpenStudy (anonymous):

oh i believe the answer is y = 1/4x + 3/2

OpenStudy (anonymous):

Okay thank you, Do you mind explaining how you got that so i know what to do for the future?

OpenStudy (anonymous):

also its negative 1/4x

OpenStudy (anonymous):

@timanti

OpenStudy (anonymous):

oh ok and you're welcome

OpenStudy (anonymous):

So its still D even though its negative?

OpenStudy (anonymous):

no its most likely y = -4x -6

OpenStudy (anonymous):

its minus 6 not negative xD

OpenStudy (anonymous):

So would it still be correct?

OpenStudy (anonymous):

yes